OpenStudy (anonymous):

Daniela has $35.20 in dimes and quarters. She has 2 more dimes than quarters. How many quarters does she have?

OpenStudy (anonymous):

let x be dimes and y be quarters 0.1x+0.25y=35.2 [1] x=y+2 Sub x into [1] 0.1y+0.2+0.25y=35.2 0.35y+0.2=35.2 0.35y=35 y=100 She has 100 quarters :)

OpenStudy (anonymous):

THANKS A BUNCH!